john throws the ball straight upward and after 1 second it reaches its maximum height then it does free fall motion which takes 2 seconds. What is its maximum height of the ball?


1
Expert's answer
2021-09-26T08:39:34-0400

The ball falls from the highest point of its trajectory within 2 seconds:


"h=\frac{gt^2}{2}=\frac{9.8·2^2}{2}=19.6\text{ m}."